An object is projected from the ground with an initial velocity of 28.5m/s at 350 above the horizontal. Find the following:

(a) horizontal and vertical components of its initial velocity,

(b) time to reach the maximum height,

(c) time of flight,

(d) maximum height,

(e) the speed at the maximum height

(f) range, and

(g) velocity upon striking the ground.


1
Expert's answer
2022-05-20T04:53:28-0400

Given:

"v_0=28.5\:\rm m/s"

"\theta=35^\circ"

"g=9.8\:\rm m/s^2"


(a)

"v_{0x}=v_0\cos\theta=28.5\cos35^\circ=23.3\:\rm m/s"

"v_{0y}=v_0\sin\theta=28.5\sin35^\circ=16.3\:\rm m/s"

(b)

"t_1=\frac{v_{0y}}{g}=\frac{16.3}{9.8}=1.66\:\rm s"

(c)

"t_2=2t_1=2*1.66=3.32\:\rm s"

(d)

"h_{\max}=\frac{v_{0y}^2}{2g}=\frac{16.3^2}{2*9.8}=13.6\:\rm m"

(e)

"v=v_{0x}=23.3\:\rm m/s"

(f)

"R=v_{0x}t_2=23.3*3.32=77.4\:\rm m"

(g)

"v_{fx}=v_{0x}=23.3\:\rm m/s"

"v_{fy}=-v_{0y}=-16.3\:\rm m/s"